### v4

# Representing Molecules with Algebraic Data Types: Beyond SMILES and SELFIES

Abstract
Algebraic data types (ADTs) let a representation specify—at the type level—what molecular values are valid and what transformations are meaningful. We propose a molecular representation as a family of typed ADTs that separates (i) constitution (Dietz-style bonding systems), (ii) 3D coordinates and stereochemistry, and (iii) electronic-structure annotations. This separation makes invariants explicit, supports deterministic local edits, and provides hooks for symmetry-aware and Bayesian modeling. These data structures, allow us to consider how the representation constrains operations which may be performed over them. Types make invalid manipulations unrepresentable and make it easier to define meaningful priors/likelihoods over generative models (programs with sample and score operations). Unlike string-based formats, the ADT exposes chemical structure directly; validity conditions (e.g., valence and symmetry constraints) can be enforced by construction and checked deterministically during transformations. We optionally attach electronic-structure annotations (shell/subshell/orbital metadata) to atoms when such information is available; we do not attempt to compute orbitals in this work.
We sketch Bayesian probabilistic programming via an integration with LazyPPL, a lazy probabilistic programming library; molecules can be made instances of a group under rotation to support geometric learning settings where molecular properties are invariant under rigid motions and relabellings;; and the framework’s flexibility is demonstrated through an extension to represent chemical reactions.
We provide a Haskell library implementing the representation, released under an OSI-approved open-source license and archived with a DOI.
Scientific contribution: We introduce a typed molecular data model based on ADTs with a Haskell reference implementation. The model combines a Dietz-style constitution layer with geometry and electronic annotations, and it is designed to support validation, deterministic transformations, and integration into Bayesian and geometric learning workflows.
keywords:
Molecular Representation, Algebraic Data Types, Functional Programming, Bayesian Inference, Geometric Deep Learning, Reaction Representation
1 Introduction
The choice of molecular representation affects cheminformatics workflows such as molecular property prediction and generative design [1]. More expressive representations can make it easier to encode the information required by downstream models, but performance also depends on computational cost, ease of integration, and how reliably the representation supports valid edits, comparison, and conversion [2, 3]. Here we explore an algebraic approach: representing molecules as typed values built from constructors for atoms, connectivity, stereochemistry, and geometry, together with composable operations on that structure.
Before continuing with the article, we introduce the terms of art used throughout. These informal definitions are tuned to our setting; for formal statements see [4, 5].
Definition 1 (Data types and algebraic data types (ADTs)).
A data type is a compile-time contract specifying (i) which values exist and (ii) which operations are meaningful on those values. An algebraic data type (ADT) defines values by a finite family of constructors (products/records and sums/variants, possibly recursive) and is deconstructed by pattern matching. In this work, we represent molecules as ADTs so that parsing, validation, and transformation operate on structured molecular objects (atoms, bonds, bonding systems, geometry, and annotations) rather than on ad hoc string edits.
Using an ADT turns “chemical validity” into explicit invariants attached to constructors and validators. This supports deterministic traversal/feature extraction, localized edits that preserve well-formedness, and principled interoperation layers (import/export) that do not leak toolkit-specific internal conventions into the core representation.
Despite decades of work, string encodings such as SMILES [6] and, more recently, SELFIES [7] remain widely used interchange and model-input formats [8, 9]. Their convenience for storage and exchange comes with well-known limitations: standard SMILES is primarily graph-level and cannot directly represent many complex bonding patterns (e.g. general delocalisation or multicentre bonding); strings do not include explicit 3D information (coordinates, torsions), and stereochemical coverage is uneven across edge cases. In addition, SMILES is syntactically brittle – models can emit invalid strings—and small token edits can induce large structural changes, complicating learning and property prediction [7]. Finally, token sequences do not expose symmetry actions (e.g. $E(3)$) or typed structure directly, so geometric or Bayesian workflows typically require additional parsing and featurisation steps.
Fingerprints encode molecules into fixed-length feature vectors for retrieval, clustering, and QSAR. They are effective as descriptors, but can be lossy with respect to stereochemistry and geometry and are not, by themselves, canonical molecular representations. In this paper we therefore treat fingerprints as downstream descriptors derived from a structured representation, not as the representation itself. These approaches may be effective for retrieval and similarity ranking, but they can be lossy with respect to stereochemistry or geometry, and they often require careful choices of similarity metrics and calibration when used for learning or uncertainty estimation [10, 11, 12]. In this paper we treat fingerprints as downstream descriptors rather than as a primary molecular representation.
Krenn et al. also proposed a “molecular programming language”, noting that strings can encode Turing‑complete languages [1]. That observation is correct, but not prescriptive. Rather than extending a string grammar until it behaves like a language, one can treat the molecule itself as a first‑class program object in an ordinary programming language, with explicit types and operations defined on it. In a functional paradigm, ADTs paired with pure functions may provide composability, equational reasoning about transformations on molecules, and immutability, helping to make invariants and exceptions explicit. We propose that if molecules are to be generated, validated, optimised and transformed by programs, representing them directly as typed data – rather than as tokens that must be decoded – offers a more transparent and verifiable route.
We first review limitations of common representations (string encodings, fingerprints, and widely used in-memory graph objects) with a focus on expressiveness, edit locality, validation, and compatibility with geometric and Bayesian modeling [2, 3]. We then introduce our ADT-based representation and its Haskell implementation, including (i) a Dietz-style constitution layer, (ii) a coordinate layer, and (iii) electronic annotations. Finally, we provide examples (e.g., benzene) and a prototype probabilistic-programming integration to illustrate how structured molecular values can be used directly in modeling code.
A brief terminological note: conflating a storage or transmission format, e.g. whether it is transmitted as a string on Wikipedia (e.g., a SMILES or InChI string) or as an SDF file with a computational data type (the in‑memory structure and its operations) has led to methodological confusion. We argue that this conflation has nudged cheminformaticians toward grammar engineering and away from semantically faithful, geometry‑aware modeling. The algebraic view places data types first: molecules as typed values with constructors for atoms, bonds, stereochemistry, and 3D information (including dihedral angles), supporting modelling choices where invariances, validity, and explicit structure matter (e.g., geometric deep learning and uncertainty-aware inference).
Other work has linked functional programming and cheminformatics before [13, 14], though this work was primarily focused on implementing existing cheminformatics workflows within functional programming languages, and the benefits of immutable state and type‑safety. Instead, we provide a molecular representation and the surrounding context for the importance of considering data types when choosing a representation.
This work presents a reference implementation rather than a final standard. Our goal is to demonstrate how an ADT-based approach can make representational assumptions explicit, support composable transformations, and provide a foundation for future evaluation and extensions in cheminformatics.
1.1 Molecular Graphs and the Dietz representation
Multigraphs retain familiar binary bonds while explicitly encoding delocalised, multi-centre electron sharing without proliferating ad hoc bond types [15].
1.1.1 Molecular Graphs
Many molecular representations are built on molecular graphs and thus inherit their strengths and limitations [16]. String encodings such as SMILES and SELFIES decode to molecular graphs, and SDF molfiles organise “blocks” for atoms and bonds.
Definition 2 (Graph).
A graph $G=(V,E)$ is a pair consisting of a set of vertices $V$ and an edge set $E\subseteq V\times V$. [^1] Set-theoretic notation used once: $V$ is the vertex set and $E\subseteq V\times V$ the set of ordered pairs $(u,v)$; for undirected graphs $(u,v)$ and $(v,u)$ are treated equivalently.
A molecular graph has vertices as atoms and edges as bonds; such graphs are connected, undirected, and labeled. Vertex labels may include element and charge; edge labels may include bond type/order or stereochemical annotations [15, 16].
Graph representations let models exploit chemical graph theory directly. However, because edges relate two vertices only, simple graph encodings can be awkward for delocalized or multicenter bonding patterns and for representations where bond order is not well-defined without additional electronic-structure context [16, 17]. Edge labels (e.g., rational bond orders) do not capture how the same electrons contribute to multiple bonds, and encoding delocalisation by multiplying bond types becomes ad hoc. As Dietz noted:
“Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution.”
A single graph also cannot represent tautomers, typically requiring multiple structures. Hypergraph and multigraph models have been proposed to address some of these issues [18], but have seen limited use [16].
1.1.2 Molecular Hypergraphs
A hypergraph allows edges (hyperedges) to connect any number of vertices. In molecular hypergraphs, 2-vertex hyperedges represent localised bonds; hyperedges with more than two vertices represent delocalised electron systems and can be labeled by the number of shared electrons [15].
Dietz cautioned that multi-atom hyperedges erase binary neighborhood information:
“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal concerning the way in which the electrons are delocalized over these atoms” [15].
Because multi-vertex hyperedges suppress pairwise adjacencies, they hinder algorithms that rely on binary neighborhood structure, and Dietz therefore rejects hypergraphs for constitutional representation [15].
1.1.3 Multigraphs
A multigraph $G=(V,E)$ has vertex set $V$ and a multiset (bag) $E$ of edges—edges may repeat. Multigraphs appear naturally when multiple bonding systems contribute between the same atom pair.
An edge in the multigraph corresponds to a bonding relation between a specific atom pair $\{i,j\}\in F$ and is labeled by the identifier of its bonding system and the shared electron count $s$. Localised bonds are the special case $|F|=1$; delocalised or multi-centre bonding is captured by a single bonding system whose $F$ contains multiple atom pairs (e.g., the benzene $\pi$ sextet spans six C–C pairs). Multiple bonding systems may contribute between the same two atoms (e.g., $\sigma$ and $\pi$ components), which is why a multigraph—rather than a simple graph—is required [15].
1.1.4 Dietz representation
The Dietz representation provides a computational framework capable of representing any chemical structure that can be described by a structural formula, overcoming the limitations of (non-multi)graph-based and hypergraph-based methods (including string encodings that decode to molecular graphs) [15]. Dietz factors structure into constitution, configuration, and conformation. Here we focus on constitution, since our ADT supplies 3D coordinates separately.
Definition 3 (Constitution).
The constitution of a molecule is an ordered pair $C=(V,B)$ where $V$ is the set of atoms and $B$ the set of bonding systems:
$V=\{\,(u,j,A)\mid u\in\mathbb{N}_{0},\ j\in\mathbb{Z}_{+},\ A\in\Sigma\,\},$
where $(u,j,A)$ represents an atom with $u$ unshared valence electrons, unique index $j$, and atomic symbol $A\in\Sigma=\{H,C,O,\dots\}$. Let $J=\{\,j\mid\exists\,u,A:\ (u,j,A)\in V\,\}$ and $\binom{J}{2}=\{\,\{i,k\}\subset J\mid i\neq k\,\}$. Each bonding system is a pair
$B\subseteq\{\,(s,F)\mid s\in\mathbb{N}_{0},\ \emptyset\neq F\subseteq\tbinom{J}{2}\,\}$,
where $s$ is the number of shared electrons in the system and $F$ is a nonempty set of unordered pairs of vertex indices (atom pairs) over which those electrons are distributed.
Examples.
Localised bonding: a 2-electron covalent bond between atoms $i$ and $j$ is $(2,\{\{i,j\}\})$. Delocalised bonding uses $|F|>1$, e.g., benzene’s $\pi$ sextet:
$(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\}).$
Example 1 (Constitution of Hydrogen)
$V(\mathrm{H}_{2})=\{(0,1,H),(0,2,H)\},\qquad B(\mathrm{H}_{2})=\{(2,\{\{1,2\}\})\}.$
Example 2 (Constitution of Benzene)
[…]
Derived quantities (Dietz).
Let $v(x)$ be the valence of the isolated atom, $u(x)$ the unshared electrons, $B_{x}$ the bonding systems involving atom $x$, $n(b)$ the electrons in $b$, $p(b)=|F_{b}|$ the number of atom pairs in $b$, and $p_{x}(b)$ the number of pairs in $b$ that contain $x$. Then the electrons belonging to $x$ in the structure are
$v_{s}(x)\;=\;u(x)\;+\;\sum_{b\in B_{x}}\frac{n(b)\,p_{x}(b)}{2\,p(b)}\,$,
and the formal charge is $c_{s}(x)=v(x)-v_{s}(x)$. The formal bond order between $x$ and $y$ is
$\mathrm{bo}(x,y)\;=\;\tfrac{1}{2}\sum_{b\in B_{\{x,y\}}}\frac{n(b)}{p(b)}\,$,
with $B_{\{x,y\}}$ those bonding systems for which $\{x,y\}\in F_{b}$ [15]. (Electrostatic interactions (e.g., ionic or hydrogen-bond contacts) can be modeled, if desired, as bonding systems with s = 0 to record neighborhood relationships without assigning shared-electron counts.)
Induced labeled multigraph (from $C=(V,B)$).
Let $E^{\ast}=\{\,(\{i,j\},b)\mid b=(s,F)\in B,\ \{i,j\}\in F\,\}$. The induced multigraph is $G^{\ast}=(V,E^{\ast})$ with edge-labeling $\ell\colon(\{i,j\},b)\mapsto(\operatorname{id}(b),s)$. This makes explicit how $(V,B)$ induces a labeled molecular multigraph while preserving pairwise adjacencies and capturing delocalisation via edge-induced subgraphs labeled by bonding-system identifiers.
Benzene can be visualised as Lewis Structure as follows:
Figure 1: Benzene – Kekulé structure on the left and on the right Benzene’s more general form. Image Credit: Oliver Goldstein (Author).
The delocalised $\pi$ electrons of the benzene ring are described in a single bonding system, as the final element of the set, and separately from all other covalent bonds in the structure.
Equivalently, the constitution can be viewed as a multigraph together with metadata grouping edges into electron-sharing systems.
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original article. The lack of labeling requirements for delocalised bonding stands in contrast to string-based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
[…]
$\quad(6,\{\{1,2\},\dots,\{1,6\},\{2,3\},\dots,\{6,2\}\}),$
[…]
Figure 2: Three‑dimensional view of ferrocene showing the metallocene “sandwich” geometry. Image Credit: [19].
[…]
Despite satisfying the three requirements identified in [1] for representation beyond organic chemistry—namely, the ability to represent delocalized bonding, explicit labeling of terminal hydrogen atoms, and advanced stereochemistry such as in cumulenes—Krenn et al. describe the Dietz representation (apparently mischaracterized as using hypergraphs) as one that “leads to complicated nested sets of brackets that may be hard to comprehend” [1].
Although set-theoretic notation may be unfamiliar, the definition or “grammar” of the constitution of a molecule can be efficiently described. This stands in contrast to SMILES strings or SELFIES strings (which as an additional challenge contains overloading of symbols in the grammar [1]).
1.2 Strings
1.2.1 SMILES
SMILES encodes molecular graphs as ASCII strings using atoms, bonds, branches, and ring closures, with optional isomeric annotations for stereochemistry. As a line notation, SMILES is compact, widely used [16, 17, 1, 20, 21, 22, 23] and widely interoperable; however, it does not store three‑dimensional coordinates or conformers and, like any graph‑level encoding, it must rely on external geometry to recover dihedrals and conformations.[24] In practice, limitations arise around (i) resonance/tautomerism and delocalisation (which require multiple structures or conventions), (ii) the breadth of stereochemical cases that can be expressed at the graph level, and (iii) variation across toolchains (e.g., aromaticity perception, canonicalisation).[17] These are active areas of improvement: OpenSMILES provides a community spec, and newer proposals such as BALSA offer a formally specified subset with explicit syntax/semantics to reduce ambiguity across implementations [25].
While SMILES can represent aromaticity through specific syntax[6], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
The encoding of stereochemistry in SMILES has notable constraints. While tetrahedral centers are well-supported through isomeric SMILES, a common extension, the encoding of more complex stereochemical features, such as dynamic or rotational stereoisomerism seen in atropisomers, remains unsupported. Isomeric SMILES can represent certain forms of axial chirality in cumulenes and handle non-tetrahedral stereocenters, including allene-like, trigonal-bipyramidal, and octahedral configurations [24]. However, these representations rely on 2-dimensional graph-based encoding and lack support for nuanced 3D spatial arrangements required for complex organometallic systems or multi-centre bonding scenarios. The absence of three-dimensional (3D) spatial data further limits SMILES’ utility for applications requiring dihedral angles or precise molecular conformations, a critical aspect of protein-ligand modeling and other machine learning tasks such as property prediction [26, 27, 28].
Beyond these limitations, SMILES suffers from unique structural and functional shortcomings such as syntactic invalidity [7, 16, 17], non-local encoding of features, and non-unique representations of molecules, all of which can impair machine learning models as the models must implicitly learn the SMILES syntax, which can be difficult, and require larger training sets and wasted computation [16, 29].
“Syntactic invalidity” means that not all valid SMILES strings correspond to valid molecules. In Machine learning contexts, algorithms may output syntactic nonsense, and must re-sample or require larger training sets to ensure robustness [16, 30, 7]. Even though one paper claims that invalid SMILES should be seen as beneficial to chemical language models [31], it should be noted that the perceived improvement is only relative to comparisons between SMILES and SELFIES models, and likely reflects the models’ enhanced ability to infer SMILES syntax when generating valid molecules.
In SMILES, small changes in the string can lead to drastic changes in molecular structure, and vice versa – adjacent atoms in the molecule can be far apart in the string [29]. As well as impacting readability, the non-local encoding of structural information can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules [29].
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [6]. As well as complicating database searches [17], machine learning algorithms must discern the equality of different representations. Although some machine learning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets, and have claimed improvements to their models as a result [32, 33], these gains may reflect regularisation effects from representing the same graph with multiple strings, but they also underline that the learning objective must spend capacity on the string grammar rather than chemistry. This motivates representations whose locality and invariants are explicit.. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
SMILES does not explicitly represent molecular symmetries (graph automorphisms) as first-class objects. For example, the Cyclopentadienyl anion in Figure 3 can be written as [cH-]1cccc1, but the rotational symmetry evident in Fig. 1 is not encoded in the string. Moreover, because SMILES is non-unique, alternative strings such as 1cccc1[cH-] represent the same structure without any explicit indication of equivalence at the representation level.
Figure 3: The Cyclopentadienyl anion. Although the structure displays clear rotational symmetry, this is not indicated by the SMILES representation: [cH-]1cccc1. Image Credit: Oliver Goldstein (Author).
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, have led to inconsistencies between implementations and research teams [16], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
Importantly, as an untyped data format, a string does not by itself enforce semantic invariants (e.g. valence constraints, charge consistency, or well-formed stereochemical annotations); those guarantees live in external parsers, validators, and toolchain conventions.
While SMILES and its extensions remain useful tools for encoding and sharing molecular structures, its limitations constrain its applicability in modern cheminformatics and machine learning workflows. Extensions to SMILES to fix problems, and the problems they themselves introduce, are discussed in a later section.
1.2.2 SELFIES
SELFIES (Self-Referencing Embedded Strings) [7] addresses a major practical issue with SMILES: syntactic invalidity. They create a “100% robust” representation, which would translate to soundness in the terminology of logic or computer science. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecular graphs (soundness) and all molecular graphs can be represented (completeness – but only up to the limited CFG they have created, see Fig. 1). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [34].
In attempting to retain semantic as well as syntactic validity after mutation, SELFIES uses a table of derivation rules, with overloading of terms to encode valence information [1]. Although successful in maintaining semantic validity after mutation (with the given set of rules, for small biomolecules), and as successfully demonstrated with GANs and VAEs[7], the overloading of terms makes the grammar more complicated. The authors claim that SELFIES may be more readable than SMILES for large molecules, although they acknowledge that “Read-ability is in the eye of the beholder”[7]. In practice, SELFIES tokens can be less directly interpretable than SMILES substrings, because the mapping from tokens to familiar chemical motifs is less transparent; whether this matters depends on the use case (human inspection vs model robustness).
For machine learning, SELFIES still presents a token sequence governed by a formal grammar. Models therefore must learn token-level regularities to generate valid structures, even if the decoder guarantees validity for any terminal sequence [16]. Claims of improved “machine readability” should be interpreted cautiously: whether robustness yields systematic gains depends on the task, architecture, and evaluation protocol.
In light of earlier discussions of line-notation design trade-offs [6], the robustness – interpretability balance in SELFIES is best viewed as a pragmatic engineering choice: additional rule structure can improve validity under mutation but may reduce human interpretability compared with simpler notations.
State $\epsilon$ [F] [=O] [#N] [O] [N] [=N] [C] [=C] [#C] [Branch1] [Branch2] [Branch3] [Ring]
$X_{0}$ $X_{0}$ F 0 $X_{2}$ N $X_{3}$ 0 $X_{2}$ $NX_{3}$ $CX_{4}$ $CX_{4}$ $CX_{4}$ ign $X_{0}$ ign $X_{0}$
$X_{1}$ $\epsilon$ F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ ign $X_{1}$ ign $X_{1}$ R(N) R(N)
$X_{2}$ $\epsilon$ F =O $NX_{2}$ 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ B(N, $X_{5}$) $X_{1}$ B(N, $X_{5}$) $X_{1}$ R(N) $X_{1}$ R(N) $X_{1}$
[…]
$X_{5}$ C F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ $X_{5}$ $X_{5}$ $X_{5}$ $X_{5}$
$X_{6}$ C F =O $NX_{2}$ 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$
$X_{7}$ C F =O #N 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$ #C $X_{1}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$
NN 1 2 3 4 5 6 7 8 9 10 11 12 13 14
Table 1: Derivation rules of SELFIES [7] for molecules in the QM9 dataset.
Again, like SMILES, resonant structures are not directly expressible in SELFIES, and delocalized bonding (other than aromaticity) cannot be expressed, along with tautomers.
Since publication, SELFIES has added the ability to represent chirality, through importation of symbols from SMILES, although some forms produce errors in decoding. However, three dimensional information is not included, and it is unclear how to extend the grammar whilst retaining its simplicity and soundness. The spatial arrangement of ligands in square-planar stereochemistry, like that of cisplatin and transplatin, cannot be fully captured without 3D information, and so in exploring the chemical space, SELFIES remains unable to represent potentially crucial information for novel drug discovery and property prediction, despite being sound.
Furthermore, E-Z stereochemistry does not in general provide accurate dihedral or torsional angles. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks. As Dayalan explains ([28]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [35]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
1.2.3 Extensions
The future of SELFIES and Future Projects
In “SELFIES and the Future of Molecular String Representations” Krenn et al. (2022) [1], review the limitations of string-based molecular representations in cheminformatics and propose future directions, primarily centered around extending string-based methods.
They highlight several representational difficulties with current molecular string formats, including their inability to effectively capture macromolecular structures, crystal lattices, complex bonding (e.g., organometallics), advanced stereochemistry, precise conformations, and non-covalent interactions such as hydrogen and ionic bonding. These limitations, they argue, are not only present in SMILES and SELFIES but are features of most modern, digital molecular representations in use.
The authors also identify challenges specific to machine learning applications, such as issues arising from the non-local encoding of molecular features within strings, and the impact this has on the latent space of VAEs, as well as the impact overloading of symbols has the efficiency of generative models, as well as human readability.
Macromolecules and crystals.
For macromolecules, Krenn et al. discuss existing approaches such as CurlySMILES, BigSMILES, and HELM, each of which introduces distinct syntactic modifications to account (or not, as the case may be) for repeating subunits, complex connectivity, and stochastic relationships between subunits. They propose BigSELFIES and HELM-SELFIES as future extensions to enable the use of generative models which take strings as input for macromolecular synthesis and design.
In the case of crystals, the authors propose yet another syntax, Crystal-SELFIES, based on the labeled quotient graph of the crystal structure. The formalism of the approach is largely unexplored in the paper, however relies on representation of subunits, in a related manner as with macromolecules.
Extensions of SMILES.
SMILES remains the dominant line notation, but fixes and extensions continue to appear. DeepSMILES reduces bracket/closure errors yet does not eliminate syntactic invalidity altogether [36]. Daylight’s ecosystem introduced related languages (e.g., SMARTS/SMIRKS for queries and transforms; CHUCKLES/CHORTLES for mixtures), and OpenSMILES provides a community specification [37]. For polymers, BigSMILES extends string notation to stochastic connectivity [16]. For reactions, SMIRKS captures structural transforms but leaves conditions to external metadata; later, RInChI added a layered, direction‑aware identifier, and ProcAuxInfo was proposed for process data (yields, temperature, concentrations) [38, 17, 39]. These ad-hoc improvements continue to this day [40]. Finally, Balsa offers a fully specified, machine‑readable subset of SMILES to reduce ambiguity across implementations [25]. These innovations improve coverage but also fragment the landscape; our focus is on representations that are precise, interoperable, and conducive to learning.
Molecular fingerprints
Fingerprints encode features of a molecule into arrays for retrieval and modeling; we treat them here as descriptors, not as canonical molecular representations [10, 16]. Classical examples include atom pairs and topological torsions (topology), ECFP/FCFP (local neighborhoods), and pharmacophore pairs/triplets (functional interactions); string‑based variants operate directly on SMILES (e.g., LINGO; MAP4) [11, 10]. In practice, analysis relies on set/bit similarities rather than vector‑space algebra; the Tanimoto (Jaccard) index is standard [12]. Limitations are well known: tautomers may hash to different bitmaps [41], and hashing plus high dimensionality can induce bit collisions and sparsity that limit expressivity [29].
1.3 Existing Programming language representations:
1.3.1 Molecules and programming languages
Inductive Logic Programming (ILP) has represented molecules, notably in Srinivasan et al. [42], but the emphasis was rule induction on graphs rather than ILP as a general molecular representation, which likely limited uptake and saw it omitted from recent surveys such as Wigh et al. [17]. In this line, molecules are encoded as Prolog rules in a first-order declarative language; yet Prolog’s narrow primitive types and operational features, including negation as failure and side-effecting built-ins like read, get, and assert, mean it is not purely declarative and can complicate reasoning. Srinivasan et al. used Progol, an ILP extension, to induce rules [42] (distinct from ProbLog, a probabilistic logic programming language). More broadly, de Meent et al. argue modern AI progress stems from tools (e.g., NumPy) that automate gradients, extended by probabilistic programming to generic modeling with uncertainty [43]; early ILP work did not integrate such machine-learning techniques [42].
1.3.2 Functional languages and cheminformatics
Work using functional programming in cheminformatics has been done before, however the focus of such efforts were usually related to ease of writing or safety of code, or efficiency of functional programming languages, rather than attempting to introduce a new representation by explicitly considering data types.
Ouch [44] is a chemical informatics toolkit written entirely in Haskell, released in 2010. It was released “in the hope that it will be useful”, but has seen little development since 2013. Internally, molecules are represented with a molecular graph, implemented as an algebraic data type, where bond types are categorically labeled (e.g. “sigma”, “pi”, or “aromatic”), and without any support for stereochemical features or 3D information.
Chemf[14] is a toolkit written in Scala. Its purpose was not to create a molecular representation, but to permit existing cheminformatics workflows to utilise the functional programming paradigm, and to avoid the problems which can occur with mutable state and null pointers. The representation used is a 2D molecular graph written as a connectivity list.
“Radium” [45] is a simple Haskell library. It has support for atomic orbital information, but molecules may only be encoded as a linear notation based upon SMILES strings. Interestingly, the data types used for molecular representation are algebraic, and while being based upon SMILES syntax, are not in fact “strings”.
Other work [13] has suggested using OCaml for cheminformatics, however this work was primarily written as an introduction to functional programming for chemists, with a granular explanation of specific functions, and a focus on ease of writing code, type safety, and efficiency, such as parallelization.
We are not aware of prior work that advises cheminformaticians use constructs rooted in functional programming languages directly in the representation of molecules themselves. Additionally, unlike previous work in functional programming, we attempt to expand the scope of molecular representation, and explicitly consider the role of the data type and how this affects machine learning, including with concepts such as Bayesian machine learning and Geometric deep learning.
1.3.3 RDKit as a Representation
RDKit is a mature and widely adopted toolkit for parsing, editing, and interoperating molecular data. Our aim, however, is different: we seek a foundational representation suited to differentiable, symmetry-aware Bayesian machine learning, where invariants are explicit and transformations are compositional. Structurally, RDKit’s core abstraction is a mutable, label-based 2D molecular graph, with three-dimensional geometry stored separately in Conformer objects and stereochemistry or delocalisation encoded via enumerated labels rather than algebraic structure [46]. This separation complicates formal reasoning about global invariants (e.g., consistency between topology, geometry, and electronic interpretation) and makes it difficult to treat molecules as immutable mathematical objects. In contrast, we adopt an algebraic, typed substrate in which constitution, geometry, and electronic annotations are integrated into a single structured value, enabling deterministic traversal, principled validation, and compositional transformations. RDKit therefore plays a complementary role in our workflow—as an interoperation and import/export layer—rather than serving as the semantic core of the representation.
2 Methodology
2.1 Our Algebraic Data Type
A typed molecular representation.
We implement a molecular Algebraic Data Type (ADT) in Haskell [47]—a lazy, purely (albeit potentially non‑terminating and exception‑raising) functional, type‑safe language whose algebraic data types are effectively isorecursive, exposing an implicit fold/unfold discipline via constructors and pattern matching (e.g., data List a = Nil | Cons a (List a)). Two language features are central to our design. First, type classes provide named bundles of operations that a type implementing such typeclass promises to support, i.e. instances supply the implementations; this yields principled operator overloading and generic code [48]. We use type classes to give uniform APIs for pretty‑printing, equality, hashing, and distance/selection queries over molecular parts. Second, static typing checks these contracts before execution. In the Hindley–Milner tradition, the compiler infers principal (most general) types automatically, minimizing annotation burden, while type safety (progress+preservation) ensures well‑typed programs do not “get stuck” [49, 4, 50]. We exploit this to encode chemical invariants in types and to reject ill‑formed constructions at compile time.
Constitution (Dietz valence‑multigraph).
At the core, we adopt the Dietz valence‑multigraph for molecular constitution [15]. By construction, it supports delocalised and multicentre bonding, organometallic and electron‑deficient systems, resonance alternatives, fractional/zero bond orders, ionic bonding, and explicit hydrogens. This generality addresses three limitations identified for extending SELFIES beyond organic chemistry—delocalisation, explicit hydrogen counts, and complex stereochemistry [1]. The constitution layer exposes class‑constrained constructors and pattern matches: instances capture admissible compositions (e.g., what counts as a bond or atom in a given sublanguage), while static typing rules out structurally impossible assemblies before runtime.
Coordinates
A coordinate layer attaches three‑dimensional atomic positions to the constitutional graph. When present, stereoisomers are distinguished and geometric features (bond lengths, angles, torsions) are computable with modest additional storage. The layer is strictly separated from constitution to enable geometry‑free symbolic manipulation. Type‑class‑based interfaces present the same traversal and query operations to both geometry‑free and geometry‑aware values, while static types make the presence/absence of coordinates explicit to clients of the API.
Electronic structure (Orbital ADT)
We implement a dedicated Orbital ADT to model shells, subshells, and orbitals with occupancy. Conceptually, it factors into (i) principal shell, (ii) subshell type (s, p, d, f, …), (iii) orbital index within a subshell, and (iv) electron population (integer or fractional). The Orbital ADT can be associated with atoms or bonds, enabling explicit representation of electron counts, lone pairs, and multicentre electron sharing. Because Orbital is first‑class, the same fold/unfold discipline applies. Here, type classes provide the admissible operations (e.g., population queries, spin projections), while instance‑guided smart constructors and static types constrain construction to chemically valid states (e.g., occupancy ranges), catching errors early.
Type‑driven composition and traversal.
Haskell type classes specify admissible molecular components and compositions, constraining construction of ill‑typed structures. Standard higher‑order and monadic combinators (e.g., map, fold, forM_) then provide idiomatic iteration, filtering, and stateful exploration over molecules, coordinates, and orbitals. This combination—type‑class abstraction plus static typing and inference—yields reusable, strongly typed building blocks for downstream algorithms without sacrificing genericity.
Reactions and probabilistic programs.
The same typed representation extends to elementary reactions by lifting the constitution/coordinate/orbital layers over multisets of molecules with atom‑mapping annotations. Type classes define the reaction‑level interfaces (e.g., how to apply an atom map, how to aggregate stoichiometry), and static typing ensures that these lifts preserve invariants across reactants and products. To illustrate compatibility with Bayesian machine learning, we provide a reference implementation of a probabilistic model in Lazy PPL that consumes/produces values of the ADT. These implementations are architectural: they define data, interfaces, and semantics used by our experiments in later sections; evaluation results are reported separately.
Practical role.
Beyond in‑memory computation, the ADT serves as a stable, serialisable format for storage, retrieval, and exchange of molecular information. Because molecules are constructed directly as well‑typed programs in the ADT’s grammar, learning and inference can operate on the representation itself—without ad hoc encodings or external context. Type classes deliver uniform tooling (pretty‑printing, hashing, equality), while static typing and inference deliver early error detection and refactor‑friendly guarantees that are essential for reliable cheminformatics pipelines.
2.2 Haskell
In Haskell, data defines algebraic data types (ADTs) as closed sets of constructors, so every value has a known shape and functions can pattern‑match exhaustively. Type classes declare which operations a type supports—together with intended laws that instances should satisfy—enabling ad‑hoc polymorphism (e.g., Eq for ==). Optionality is explicit via Maybe/Nothing (e.g., a missing SubShell), avoiding nulls and exception‑driven control flow. Haskell’s static types and purity shift many failures to compile time, which is valuable for scientific ML pipelines.
Our contribution is to bring these guarantees to probabilistic programming for cheminformatics. Existing Haskell Probabilistic Programming Languages (PPLs) (e.g., LazyPPL) support concise Bayesian models where lazy evaluation defers costly sampling/simulation until needed—useful for molecular property modeling and design [51, 52, 53, 54, 55, 56]. The ADT layer integrates naturally with robust graph/string encodings used in de novo design, such as SELFIES and DeepSMILES [1, 36]. (ADTs and purity are not unique to Haskell; Haskell simply offers a mature, succinct realization.)
2.3 Record Syntax
In our reference implementation, atoms are typed records (Atom{atomID, atomicAttr, coordinate, shells}). Named fields give local, direct access to exactly the features needed for ML or simulation (e.g., symbol (atomicAttr a), x (coordinate a)), avoiding whole‑molecule string parsing and touching only the required data – lighter than non‑local encodings such as SMILES/SELFIES [1, 36]. The model can evolve by adding fields (e.g., charges, isotopes) with minimal impact when code uses named fields.
1updateXCoordinate :: Double -> Atom -> Atom
2 updateXCoordinate newX a =
3 a { coordinate = fmap (\(Coordinate _ y z) -> Coordinate newX y z) (coordinate a) }
Listing 1: Updating coordinates: structure-preserving local edit
Explanation: The pattern destructures the nested Coordinate, replaces x with newX, and reconstructs the outer Atom; other fields are passed through unchanged. This could also return a part of the structure.
3 Applications and Evaluation
3.1 The Molecule Algebraic Data Type
1newtype AtomId = AtomId Integer deriving (Eq, Ord, Show, Read)
2newtype SystemId = SystemId Int deriving (Eq, Ord, Show, Read)
3newtype NonNegative = NonNegative { getNN :: Int } deriving (Eq, Ord, Show, Read)
4
5-- Canonical undirected edge (store once; enforce i<=j in ctor)
6data Edge = Edge AtomId AtomId deriving (Eq, Ord, Show, Read)
7
8data BondingSystem = BondingSystem
9 { sharedElectrons :: NonNegative, memberEdges :: Set Edge } deriving (Eq, Show, Read)
10
11data ElementAttributes = ElementAttributes
12 { symbol :: AtomicSymbol, atomicNumber :: Int, atomicWeight :: Double } deriving (Eq, Show, Read)
13
14data Atom = Atom
15 { atomID :: AtomId, attributes :: ElementAttributes, coordinate :: Coordinate, formalCharge :: Int }
16 deriving (Eq, Show, Read)
17
18data Molecule = Molecule
19 { atoms :: Map AtomId Atom, localBonds :: Set Edge, systems :: [(SystemId, BondingSystem)] }
20 deriving (Eq, Show, Read)
Listing 2: Core molecular ADT: Dietz-style constitution with explicit atoms, $\sigma$-adjacency, and electron-pool bonding systems.
[…]
Chemically meaningful quantities are derived rather than hard‑coded: formal bond orders and per‑atom electron counts are obtained by summing each system’s fractional share over its member edges, so the record stays minimal but interpretable.
Canonical undirected edges and ordered maps/sets make the structure deterministic and easy to validate (idempotent insertions, unambiguous lookups).
These design choices follow Dietz’s rationale and examples (benzene, diborane, ferrocene), which advocate “bonding systems” spanning multiple pairs while retaining pairwise information.
An Atom records its stable identifier (atomID :: AtomId), immutable elemental metadata (attributes :: ElementAttributes), a Cartesian coordinate :: Coordinate, electronic shells :: Shells, and an explicit formalCharge :: Int. Elemental metadata comprises symbol :: AtomicSymbol, atomicNumber :: Int, and atomicWeight :: Double. Coordinates are expressed in Å via the units type Angstrom, ensuring unit safety for downstream geometry. The Shells type is re‑exported from the orbital layer to allow atoms to carry ground‑state (or user‑specified) configurations. Because localBonds is a set, adding the same edge twice is idempotent and cannot inflate degree artificially.
1newtype AtomId = AtomId Integer deriving (Eq, Ord, Show, Read)
2newtype SystemId = SystemId Int deriving (Eq, Ord, Show, Read)
3
4-- Non-negative electron counts for Dietz pools
5newtype NonNegative = NonNegative{getNN :: Int} deriving (Eq, Ord, Show, Read)
6
7-- Canonical undirected edge: the constructor enforces ordering
8data Edge = Edge AtomId AtomId deriving (Eq, Ord, Show, Read)
9
10-- One Dietz bonding system: s shared e- over a set of edges, with an optional tag
11data BondingSystem = BondingSystem
12 { sharedElectrons :: NonNegative
13 , memberAtoms :: Set AtomId
14 , memberEdges :: Set Edge
15 , tag :: Maybe String
16 } deriving (Eq, Show, Read)
Listing 3: Dietz constitution primitives used by the ADT. mkEdge canonicalises undirected pairs; BondingSystem pools $s\geq 0$ shared electrons over a set of member edges.
Listing 3 realises the Dietz constitution $C=(V,B)$ in a typed record: atoms is the vertex set $V$; localBonds holds canonical undirected $\sigma$-adjacencies; and systems is a finite family $B=\{(s,E)\}$ of bonding systems, each pooling $s\!\geq\!0$ shared electrons over a nonempty set $E$ of atom–atom edges. Pairwise neighbourhoods remain explicit (via edges) while delocalised or multicentre effects are handled uniformly by the electron pools – precisely the computer‑oriented counterpart of structural formulas advocated by Dietz.
Smart constructors and invariants. Updates pass through small “smart” constructors: (i) mkEdge canonicalises undirected pairs so set semantics are by construction (idempotent insertion, unambiguous lookup); (ii) mkBondingSystem takes a pool size $s\in\mathbb{N}_{0}$ and an edge set $E$, derives the cached atom scope $\mathrm{atoms}(E)$, and rejects ill‑formed inputs (negative $s$, empty $E$, or non‑canonical edges). Despite, potentially being overridden, maintaining the use of these constructors ensures the invariant $\mathrm{memberAtoms}=\mathrm{atoms}(\mathrm{memberEdges})$ and keeps systems a well‑formed finite family, in line with the paper’s requirement that bonding systems be defined over sets of atom pairs.
The record stores primitives; chemically useful quantities are derived: per‑atom electron use $e(v)$ and per‑edge effective order $\mathrm{order}(e)$ come from summing each system’s fractional share over its member edges, rather than hard‑coding fractional labels. Coordinates attach configuration/conformation without changing constitution. Deterministic containers (Map/Set) plus canonical edges make validation and permutation‑invariant algorithms straightforward, and the Dietz pooling formalism scales from localised bonds to general delocalisation without inventing ad‑hoc bond types.
3.2 Example: Benzene
The same molecule can be specified in a considerably more compact functional style. However, that approach introduces advanced idioms that trade clarity for brevity. Because the aim here is pedagogical transparency as well as correctness, we retain the longer, fully explicit representation in Appendix A.1.
How to read Appendix A.1 (Dietz constitution, step‑by‑step).
  1. 1.
    Name the atoms & templates. The code first fixes stable identifiers $\texttt{AtomId}\ 1\ldots 12$ and caches metadata (elementAttributes, elementShells) for C & H.
  2. 2.
    Build atoms with 3D coordinates. Records $\texttt{c1}\ldots\texttt{c6}$ and $\texttt{h7}\ldots\texttt{h12}$ are constructed with IDs, element attributes, Ångström coordinates, shells, and $\texttt{formalCharge}=0$.
  3. 3.
    Link IDs $\to$ records. atomTable is a $\texttt{Map}\ \texttt{AtomId}\!\to\!\texttt{Atom}$ assembled by successive M.insert, providing fast lookup of each atom by its symbol/ID.
  4. 4.
    Lay down the $\sigma$ framework. sigmaFramework is a Set of undirected edges (mkEdge) for the six C–C ring connections and six C–H bonds—this is the pairwise adjacency.
  5. 5.
    Add the Dietz $\pi$ system. piRingEdges repeats the ring’s C–C edges; piRingSystem = mkBondingSystem (NonNegative 6) piRingEdges creates one Dietz electron pool with $s=6$ shared electrons delocalised over those edges.
  6. 6.
    Assemble the molecule. Molecule { atoms = atomTable, localBonds = sigmaFramework, systems = [(SystemId 1, piRingSystem)] }.
[…]
3.2.1 Validator
We validate every molecule immediately after construction or import, so downstream inference and learning code never explores structurally nonsensical states. The validator enforces three structural invariants—(i) every bond endpoint must refer to an existing atom, (ii) self-bonds are disallowed, and (iii) the internal bookkeeping for undirected connectivity is symmetric—and then applies a conservative, element-wise valence bound. Concretely, each atom’s “valence usage” is computed from the local $\sigma$ adjacency together with the fractional contributions implied by any Dietz bonding pools, and the molecule is rejected if any atom exceeds the element-specific maximum (getMaxBondsSymbol). Successful validation is a no-op (the moleculeis returned unchanged); failures return a short diagnostic describing which invariant was violated. The routine runs in time linear in the number of atoms and edges.
3.3 Orbital ADT
Physics-aware data and validation (Aufbau, Hund, Pauli).
In our reference implementation, quantum information is exposed through simple ADTs rather than dependent or refinement types. Orbitals, subshells, and shells are concrete datatypes (Orbital, SubShell, Shell) with integer occupancies and orientation/hybrid terms; atoms carry a shells :: Shells field inside the molecular record (modules Orbital and Molecule; see src/Orbital.hs and src/Molecule.hs). Client code does not insert electrons via an API that enforces Aufbau/Hund/Pauli at the boundary; instead, it typically obtains ground-state occupancies via a single mapping elementShells :: AtomicSymbol -> Shells and associated element tables (elementAttributes, getMaxBondsSymbol) in Constants (see src/Constants.hs).
Accordingly, the interface provides faithful data structures and convenient constructors/defaults; the implementation supplies the empirical per-element configurations and bond limits as tables. Runtime physics checks presently cover only generic structural validity (e.g., symmetric bonds and per-element maximum total bond order) via validateMolecule in Validator (see src/Validator.hs); they do not currently re-enforce Aufbau/Hund/Pauli for arbitrary edits to electronCount. We therefore document Aufbau/Hund/Pauli as invariants and provide ground-state defaults (with room to extend to ions/exceptions), while leaving strict enforcement to construction discipline or future smart constructors [57, 58, 59].
Hybridisation in the ADT.
Hybrid orbitals are represented explicitly via hybridComponents :: Maybe [(Double, PureOrbital)], a linear combination of pure $s/p/d/f$ orbitals. We treat the coefficients as mixing amplitudes whose squares sum to 1 (checked by smart constructors), and use orientation to store the spatial axis of the resulting hybrid. This captures common hybrids (e.g., $sp^{3}$, $sp^{2}$) and supports chemically meaningful predictions such as how ss‑character modulates acidity and bond geometry (Bent’s rule and modern NBO analyses) [60, 61].
3.4 Reactions
Reaction ADT
The proposed ADT can be extended to model chemical reactions. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a tool for studying and simulating chemical processes.
Listing 4 illustrates a Reaction data type, where reactants and products are represented as lists of pairs, containing a stoichiometric coefficient (a Double) and a Molecule. The conditions field captures reaction conditions such as temperature and pressure, defined with respect to a time range. Listing 5 provides an example reaction, illustrating the formation of water from hydrogen and oxygen under specific thermodynamic conditions.
Although these examples are simple, they demonstrate the flexibility of the ADT by extending it to model reaction processes. The compositional nature and modularity of the type system ensures that additional features not implemented here, such as catalysts or thermal energy, could be straightforwardly implemented. Stoichiometric or thermodynamic constraints could be enforced through use of dependent types, although this is left as future work.
1-- Reactions or transformations between chemical species
2data Reaction = Reaction
3 { reactants :: [(Double, Molecule)]
4 , products :: [(Double, Molecule)]
5 , conditions :: [Condition]
6 , rate :: Double
7 }
8
9-- Conditions under which a reaction occurs
10data Condition = TempCondition {temperature :: Double}
11 | PressureCondition {pressure :: Double}
12
13data Times = Times { startTime :: Double, endTime :: Double }
Listing 4: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. The Reaction data type is composed using the Molecule data type
1-- Example reaction: 2H2 + O2 -> 2H2O
2exampleReaction :: Reaction
3exampleReaction = Reaction
4 { reactants = [(2.0, hydrogen), (1.0, oxygen)]
5 , products = [(2.0, water)]
6 , conditions =
7 [ TempCondition 500.0
8 , PressureCondition 1.0
9 ]
10 , rate = 0.1
11 }
Listing 5: Example reaction: formation of water from hydrogen and oxygen
In contrast, string-based representations such as reaction SMILES, SMARTS, and SMIRKS attempt to encode reactions as sequences of characters [16]. Reaction SMILES, for example, separates reactants and products using the “¿¿” operator, but does not inherently encode reaction conditions, catalysts, or rate dependencies [17]. SMARTS and SMIRKS provide pattern-matching capabilities for substructure searching and reaction templates, respectively, but lack a native mechanism for enforcing reaction validity [16].
String-based representations also inherit the limitations of the string data type itself. A lack of compositionality, and difficulties in formal reasoning make it unclear how one might extend SELFIES, for instance, to define a reaction SELFIES in which all syntactically valid strings correspond to chemically valid reactions [1]. Even if such an extension were possible, enforcing network-wide properties such as reversibility, equilibrium conditions, or pathway constraints would be nontrivial.
To generalise what is written here to adequately capture the semantics of reactions and reaction networks, it would be useful to derive an implementation off of a categorical semantics such as [62]. However, this is left as future work. Chemical reaction networks align naturally with category-theoretic formalisms, where reactions can be treated as morphisms between reactant and product sets [62]. This view has been used to model biochemical pathways, open dynamical systems, and Petri nets in systems chemistry [62], making an ADT in a typed functional language with support for categorical constructs such as Functors.
While constraints such as stoichiometric balance are not enforced in our implementation, the ADT provides a foundation upon which such constraints could be incorporated in future work through dependent types or additional verification layers.
However, the compositional nature of the ADT, its ability to be amended to incorporate reaction properties, and its alignment with category theoretic formalisms make it a natural framework cheminformatics applications.
3.5 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the article, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 6.
Parsing and construction. The SDF reader constructs the atom map from V2000 atom blocks (including coordinates and element attributes), populates localBonds from bond lines, applies formal charges from M CHG annotations, and detects common six‑membered alternating rings to build a single $\pi$ system with $s=6$ over the edges of the ring. Programmatic examples (e.g. benzene) mirror this layout—six C–C $\sigma$ edges, six C–H $\sigma$ edges, and one labelled $\pi$ system spanning the ring—so text and code paths agree on $(V,B)$.
1-- Writing the molecule to a file
2writeMoleculeToFile :: FilePath -> Molecule -> IO ()
3writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
4
5-- Reading the molecule from a file
6readMoleculeFromFile :: FilePath -> IO Molecule
7readMoleculeFromFile filePath = do
8 contents <- readFile filePath
9 return (read contents)
10
11-- Example usage
12main :: IO ()
13main = do
14 -- Write the methane molecule to a file
15 writeMoleculeToFile "methane.hs" methane
16
17 -- Read the methane molecule from the file
18 molecule <- readMoleculeFromFile "methane.hs"
19
20 -- Print the molecule read from the file
21 print molecule
Listing 6: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
Definition 4 (Serialization).
Serialization is the process of converting an in-memory data structure into a standardized, external format (e.g. a byte stream or a textual encoding) so that it can be stored or transmitted and later reconstructed in its original form.
This code demonstrates the ease of serializing and deserializing molecules to and from files in Haskell. By leveraging the Show and Read instances defined for the Molecule data type and its related types, molecules can be seamlessly converted to and from their string representations, allowing them to be stored in files and retrieved later. There are algebraic laws governing these, e.g. show and then read must be the identity function and vice versa.
This functionality provides a convenient way to save and load molecular data, and is particularly useful when working with large molecules or when molecules are required across different program runs. Although storage and transmission may take place via serialisation into a human and machine-readable string of text, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the ADT can be serialized into a text string, it is crucial to recognize that the ADT itself is not merely a string.
3.6 Probabilistic Programming
Probabilistic programming languages (PPLs) automate such inference on firm mathematical footing [63]; mature systems include Pyro [64], Infer.NET [65], and Edward [66], alongside modular, composable PPL frameworks [67, 51, 52]. Relative to conventional deep learning, Bayesian methods offer interpretability and principled treatment of multiple forms of uncertainty [68, 69], and have been applied to drug design (e.g., Bayesian neural networks for toxicity prediction [56]). Bayesian inference inverts the usual programming direction: given observed outputs, infer latent inputs and parameters that most plausibly generated them. In cheminformatics, the inputs encode molecular features under a chosen representation; inference is meaningful only when sample/score are defined for that type (e.g., for a String, a prior and likelihood over strings must be specified). A generative PPL model exposes two primitives: sample (draws from specified distributions, operationally reducible to base uniform noise) and score (weights executions by the likelihood under observed data). Inference typically uses (i) Monte Carlo (e.g., MCMC), (ii) variational optimisation, or (iii) exact calculation when conjugacy permits. In this article we use Trace Metropolis Hastings [63]. For molecular representations, PPLs naturally encode structural uncertainty and property variability.
In [70, 71], there are discussions of the importance of classifiers providing not just a single output, but two key pieces of information: a class label and an associated probability. This dual output allows for a more nuanced understanding of the classifier’s confidence in its predictions. For instance, in binary classification tasks, a classifier might assign a label (e.g., ‘spam’ or ‘not spam’) and also provide the probability of the instance belonging to that class. This probabilistic information is crucial for tasks such as ranking, where items are ordered based on the likelihood of belonging to a particular class, and for making informed decisions in applications where the costs of false positives and false negatives differ. [71] emphasizes that effective classification involves not only assigning the correct label but also accurately estimating the probability associated with that label to reflect the classifier’s certainty. In the case of SMILES, an estimate of 0.5 that it the molecule is soluble may mean exactly that or it may be saying that it is an example that is has never seen before. Without a grammar that tells one what is valid and what is invalid, it relegates the probability estimates to be ambiguous and potentially meaningless. SMILES syntactic validity is decidable by parsing against the specification; the practical issue for generative modeling is that unconstrained decoders can emit invalid strings, requiring rejection sampling, constrained decoding, or alternative representations.
-- Pseudo-Haskell; domain types & PPL ops assumed in scope; M = Data.Map.
type Params = (Double,Double,Double,Double,Double)
model :: Double -> Meas (Molecule, Params)
model obs = do
let n = 3
[…]
s <- sample (uniformD [C,N,O,H])
[x,y,z] <- replicateM 3 (sample (normal 0 1))
pure Atom{ atomID=i, atomicAttr=elementAttributes s
, coordinate=Coordinate x y z, shells=elementShells s }
let pairs = [(i,j) | i <- [1..n], j <- [i+1..n]]
bs <- fmap concat \$ forM pairs \$ \(i,j) -> do
inc <- sample (uniformD [True,False])
if not inc then pure [] else do
k <- sample (uniformD [1,2,3])
let b = Bond{ delocNum = 2*k, atomIDs = Nothing }
pure [((i,j),b),((j,i),b)]
let m = Molecule{ atoms=atoms, bonds=M.fromList bs }
[b0,b1,b2,b3,b4] <- replicateM 5 (sample (normal 0 0.1))
let s = moleculeSize m
w = moleculeWeight m
a = moleculeSurfaceArea m
bo = moleculeBondOrder m
yhat = b0 + b1*s + b2*w + b3*a + b4*bo
score (normalPdf obs 0.2 yhat)
pure (m,(b0,b1,b2,b3,b4))
main :: IO ()
[…]
mh 0.1 (model obs) >>= print . map fst . take 1000 . drop 1000
Listing 7: Trace M-H sampler that grows a molecule and scores it via a linear logP model on molecular features (jitter=0.1, burn-in=1000) [67].
Some code is excluded from Listing 7 to conserve space, but its intent is described here. The module implements a single probabilistic program that grows a candidate molecule and conditions on a target property, the observed $\log P$ (which is taken from [72]). The scalar observation is loaded by observedLogPIO (e.g., from an SDF/CSV record), and the model is run under a trace Metropolis–Hastings kernel with jitter $0.1$, discarding the first $1000$ iterations (burn-in) and retaining the next $1000$ draws.
The molecular state uses a lightweight algebraic data type. A Molecule is a record with fields atoms :: [Atom] and bonds :: M.Map (Int,Int) Bond. Each Atom carries atomID :: Int, atomicAttr :: AtomAttr, coordinate :: Coordinate (a 3D point), and shells :: Shells. A Bond stores at least delocNum :: Int (an order-derived delocalisation count) and optionally atomIDs :: Maybe (Int,Int); in our representation the bond graph itself is keyed externally by atom pairs, so atomIDs may be Nothing. The map bonds is treated as symmetric by inserting both directions $(i,j)$ and $(j,i)$ for each undirected edge.
Within the generative function model, we first sample a small scaffold of $n=3$ atoms. For each $i\in\{1,\dots,n\}$, the element symbol is drawn uniformly from [C,N,O,H]; atomic attributes elementAttributes and elementShells are looked up from the symbol; and the Cartesian coordinates are drawn independently from $\mathcal{N}(0,1)$ in each dimension. This produces Atom{atomID=i, atomicAttr, coordinate, shells} with contiguous identifiers $1..n$. We then consider every unique unordered pair $(i,j)$ with $i<j$ and flip a fair coin to decide whether a bond is present. If included, the bond order $k\in\{1,2,3\}$ is sampled uniformly and mapped to delocNum = 2k; the resulting Bond is inserted under both keys $(i,j)$ and $(j,i)$ to enforce symmetry. The atoms and bonds are combined into a Molecule value m.
Property conditioning is applied through a linear predictor over standard molecular features. From m we compute moleculeSize m, moleculeWeight m, moleculeSurfaceArea m, and moleculeBondOrder m, denoted $s,w,a,$ and $bo$. The coefficients $b_{0},\dots,b_{4}$ have independent Gaussian priors $b_{i}\sim\mathcal{N}(0,0.1^{2})$. The predicted $\log P$ is
$\hat{y}\;=\;b_{0}+b_{1}s+b_{2}w+b_{3}a+b_{4}\,bo,$
and the observed value $y$ is given a Normal likelihood $y\sim\mathcal{N}(\hat{y},0.2^{2})$. In a trace M–H implementation,
a proposal perturbs a subset of the current random choices (e.g., atom labels, local $\sigma$-edges, and selected bonding-system parameters). For readability, the probabilistic-programming sketch uses a simplified subset of the full ADT; the full representation (including bonding systems) can be sampled analogously. The jitter hyperparameter controlling the proposal scale; acceptance is computed from the usual Metropolis ratio given the prior densities and the $\log P$ likelihood.
This unified program couples structure generation with property supervision: the posterior concentrates mass on those atom/bond configurations that yield features consistent with the target $\log P$, while simultaneously quantifying uncertainty in the regression parameters. Practically, this provides a compact vehicle for inverse design—proposals that move the structure are immediately scored by the property model—so the chain preferentially explores chemically plausible neighborhoods that explain the target. The same grammar-and-score paradigm transfers naturally to synthesis planning: generative moves propose intermediates, and probabilistic scores (learned or mechanistic) guide search, a pattern that aligns with successful retrosynthesis strategies in the literature [73].
3.7 Existing barriers to a general molecular representations.
String‑first encodings suffer from far more than localized bonding, hidden hydrogens, and patchy stereochemistry. They are syntactically brittle (SMILES can be invalid) and, even when made syntactically robust (SELFIES), remain semantically limited because they still decode to labeled 2D graphs [7, 1]. They are non‑unique and tool‑dependent (divergent canonicalisation and aromaticity perception), which injects noise into indexing and learning [16]. Their traversal‑based strings are non‑local with respect to chemistry, so string distances correlate poorly with geometric or energetic similarity [29]. Critically, they omit 3D coordinates and torsions—central for binding and stability—forcing external geometry generators and breaking round‑trips [26, 27]. Binary edges with integer orders cannot express delocalisation, multi‑centre bonding, hapticity, or zero‑order interactions (e.g., diborane, ferrocene) without ad hoc labels [17, 16, 15]. Coverage for non‑tetrahedral and axial/helical/planar stereochemistry is incomplete and inconsistent (e.g., cis/trans‑platin) and often needs 3D context to disambiguate [17, 24]. Tautomers/protomers and resonance are handled outside the representation; hydrogens are frequently implicit; ions, non‑covalent interactions, and spin are out of scope [74]. Polymers, macromolecules, crystals, and periodicity fragment into bespoke grammars; reactions lack typed support for stoichiometry, conditions, and rates (pushed into side‑channels like RInChI/ProcAuxInfo) [17, 38, 39]. Overall, strings are convenient formats, not principled data types, and they misalign with Bayesian and geometric ML, which need explicit structure, 3D symmetry actions, and typed priors/likelihoods [2, 3].
Our ADT is designed to encode (and our reference implementation partially realises): Dietz-style bonding systems for delocalised/multicentre bonding; explicit atom-level annotations (e.g., charge, isotope, optional hydrogens); optional coordinates supporting stereo/conformer handling; and typed extension points for reaction objects and probabilistic modelling. Where features are sketched rather than fully implemented, we mark them explicitly as future work.
These limitations, the authors argue (and we agree), are an inherent problem with string-based representations in their ability to represent complex bonding (such as with diborane and ferrocene) and advanced stereochemical features. Cis- and trans-platin are also given as examples of molecules current string-based representations cannot represent.
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in [7]:
Thus, it should be stressed again that in d- and f-block chemistry, as well as main-group organometallic compounds, it is often impossible to assign any particular bond orders without high-level quantum chemical calculations, due to the highly de localized nature of the bonding, where electrons are often spread out over a significant number of atoms, including the metal cen ter itself, the immediately coordinated atoms, and additional ligand groups.
Despite recognizing the need for a general digital molecular representation, capable of handling all of the above issues. The authors do not reconsider the suitability of string-based models, and instead, the solutions presented primarily involve introducing new grammars, expanding syntax rules, and further overloading symbols, leading to increasingly complex and specialized representations.
“Molecular programming languages”.
In Future Project #8, Krenn et al. (2022) propose a “molecular programming language”. They reason that strings can encode the powerful computational algorithms:
“Strings can store Turing-complete programming languages: In the most general case, one can store the source code of computer programs as strings. For example, a Python file is a ‘simple string’, which is executed by the Python interpreter. Python is, of course, a Turing-complete language, which means that strings can encode the most powerful computational algorithms.”
Although program source can be stored as text, treating a raw string as a semantic representation conflates storage with meaning: correctness properties live in parsers, type checkers, and interpreters, not in the character sequence itself. In the molecular setting, we therefore aim to make semantic structure explicit in the data model rather than relying on string-level mutation.
Furthermore, if the programming language is represented by a string, the question of how functions within the language are interpreted is unanswered. If they are also strings, they would at some point need to be executed by something other than a string. If they are not strings, then it remains to be understood why the authors claim that strings alone are enough to represent a Turing-Complete language, let alone the utility of such a project.
This line of work can be read as a desire for compositional semantics and robust mutation operators. Our approach provides these by modelling molecules as typed values with explicit constructors and invariants; composition is then obtained using ordinary functions over the ADT.
Future Project #9 goes beyond molecular programming languages, and proposes a programming language which is “100% robust” by finding a syntax for a programming language that however you combine elements in the instruction set, a valid program is always produced. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure: programs themselves. The feasibility and utility of “robust programming-language” objectives are outside our scope; we focus instead on ensuring that molecular representations have explicit constructors, well-scoped operations, and clear validity conditions.
We mention the definition of a domain specific language via a definitional interpreter [75]. This perspective allows the syntax of our Molecule data type to be embedded in a host language (shallow or deep), here Haskell, paired with an environment (a mapping from variables to values). In this work, we use this framing only as a conceptual bridge: rather than treating molecular strings as “programs”, we emphasise molecules as typed values with well-scoped constructors and operations that support validation, transformation, and probabilistic modelling.
We also make progress on the future projects suggested:
  • Complicated bonds (Dietz & zero-order bonds; context for FP6). We adopt the Dietz-style constitution (multigraph of bonding systems) to capture multicenter bonding and encode “zero-order” interactions as bonds with 0 shared electrons in the ADT. Krenn et al. note: “Dietz suggested a hypergraph concept …accounting for multicenter bonding,” and dashed interactions in diborane “have been termed ‘zero-order bonds’ by Clark” [1]. This work: represented as a bonding system with sharedElectrons = 0 (a “zero-order” system) and an explicit member edge set; multicentre bonding is represented by bonding systems whose member edge set has —E—¿1.
  • FP6 — Generalization of SELFIES and automatic compilation of complex rules from data (partial). Krenn et al.: “define a robust generalization of SELFIES that incorporates molecules beyond VBs” [1]. This work: we address the representational need (beyond valence-bond assumptions) not by generalizing SELFIES, but by a typed constitution + 3D configuration ADT (Dietz-style bonding systems, explicit H, advanced stereo, coordinates). We do not attempt “automatic compilation of complex rules from data.”
  • FP2 — The effect of token overloading in generative models (design relevance only). Krenn et al.: “One important question is to understand how overloading impacts ML models” [1]. This work: our ADT has no overloaded tokens (typed constructors instead of strings), removing the issue by design; we do not perform the proposed controlled study.
  • FP7 — Graph-edit rules and metaSELFIES for reactions (related, not the proposed method). Krenn et al.: “A syntactically robust reaction representation would most likely improve the performance …” and should “conserve the number of atoms …and the total charge” [1]. This work: we provide a typed Reaction ADT where such conservation checks are natural; we do not implement metaSELFIES or a graph-edit rule DSL.
3.8 Comparison to existing tools
Table 2 situates our approach relative to widely used molecular representations and toolkits by focusing on the capabilities of the core representation, rather than on downstream algorithms. String and identifier schemes such as SMILES/OpenSMILES, SELFIES, and InChI prioritise compactness, robustness, or canonicalisation, but they do not provide first-class support for three-dimensional structure, reactions, or general delocalised and multi-centre bonding; where aromaticity or resonance is handled, it is typically via flags or normalisation conventions rather than explicit structure. File formats such as SDF/Molfile can encode 2D/3D coordinates and stereochemistry, but function primarily as interchange containers rather than as editable, invariant-preserving internal models. Graph-based toolkits like RDKit and Open Babel offer rich chemistry operations, reaction handling, and broad interoperability, yet their internal representations still largely rely on conventional bond models with special-case treatments for delocalisation. In contrast, this work adopts a typed algebraic data type as the primary molecular representation, enabling explicit, first-class treatment of general delocalised bonding alongside 3D geometry, stereochemistry, and reactions, while remaining open-source and suitable as a reference model rather than merely a file format or algorithmic library.
System Core rep. General deloc. 3D Stereo Rxns OSI Notes
SMILES / OpenSMILES String Line notation for molecules;
SELFIES String Robust string representation (decoder guarantees validity under its grammar);
InChI Identifier Canonical identifier for lookup and deduplication rather than editing;
SDF / Molfile File format Exchange format: can carry 2D/3D coordinates and stereo bond annotations;
RDKit Graph toolkit Mature cheminformatics toolkit with broad interop;
Open Babel Graph toolkit Broad format conversion and chemistry operations;
This work Typed ADT Our reference implementation
Table 2: High-level comparison of molecular representations and toolkits. Here “General deloc.” means first-class support for general delocalised/multi-centre bonding (beyond aromaticity flags / kekulisation conventions). Symbols: ✓ supported; ✗ not supported; “–” not applicable.
3.9 Limitations and Future Work
We see three concrete next steps.
  1. 1.
    Coverage extensions. Add explicit support for tautomer sets, ionic/non-covalent contacts, and polymer/repeat-unit abstractions.
  2. 2.
    Interoperability. Extend parsers and round-trip tests for SMILES/InChI and reaction formats (SMIRKS/RXN) so that comparisons can be made on shared corpora.
  3. 3.
    Empirical evaluation. Benchmark parsing/validation speed and memory, and assess downstream impact on a small set of standard tasks (e.g., property prediction with symmetry-aware models) where representation choice is known to matter.
The statistical and computational efficiency of our reference implementation has not been optimised for. Our current prototype prioritises clarity over compactness. In practice, the ADT is intended as an in-memory semantic representation; serialisation can target standard compact encodings (SDF/SMILES/InChI) or a dedicated binary format. We therefore treat file-size considerations as an engineering concern rather than a representational limitation. Furthermore, by storing a molecule as function or properties as higher-order functions, it may be possible to significantly reduce the storage size of a molecule.
3.10 Future work
The ADT so far presented has the ability to be improved upon, extended, and used in empirical experiments to verify and test the representation’s utility, ease-of-use and verifiability across a number of domains in cheminformatics. Parsers from other formats e.g. SMILES strings to this ADT could also be used. Currently only parsers for .SDF files have been written.
3.10.1 Geometric Deep Learning on Molecular Representations
Future work could consist of using symmetry-aware with Geometric Deep Learning, to reduce estimation error (requiring less data to infer parameters). A group captures the set of symmetrical operations over the representation. For molecules the key properties are (i) permutations of atom IDs and (ii) rigid 3D motions (rotations/translations). Randomized SMILES already tries to mimic (i) by enumerating alternative traversals; GDL would make this automatic—any relabeling or rigid motion is handled identically (invariant for scalar properties, equivariant for vector fields). Practically, we can encode this contract in code by introducing a small Haskell Group typeclass, with instances for atom-index permutations and rigid motions; models then declare which actions they respect, and the type system enforces symmetry-safe use. This process of learning leverages algebraic structure in the domain, to reduce estimation, model and approximation error amongst other things.
3.10.2 Extending and Evaluating the ADT
The abstract data type (ADT) presented here can be extended and empirically evaluated to test its utility across diverse tasks in cheminformatics. Two complementary lines of work are: (i) improving interoperability through additional parsers, and (ii) augmenting the core representation to cover chemical phenomena not yet encoded.
Interoperability and parsers.
To broaden adoption, it is useful to support import from multiple established formats (e.g., SMILES) into the ADT; at present, we provide parsers only for SDF files. Extending the parsing layer will permit side-by-side evaluations and reduce friction when integrating the ADT into existing workflows.
Augmenting the representation
To increase expressiveness, the ADT should be extended to represent tautomerism, ionic bonding, polymerism, and selected quantum information, such as spin, which are currently not captured.
Tautomerism.
Tautomers occur in many therapeutics, from sildenafil (Viagra) to warfarin, remdesivir (used for COVID-19), tetracyclines, and other antibiotics. Explicit handling of tautomeric states can expand the reachable chemical space and improve tasks such as ligand–protein binding prediction [76, 77, 78].
Ionic bonding.
Ionic interactions are central to the stability and function of nucleic acids, proteins, and membranes, and to processes such as catalysis and ion transport [74, 79, 80]. Encoding ions and their bonding explicitly would enable modeling of charged species and electrochemical properties, supporting the design of materials including ionic liquids and solid electrolytes for energy storage and green-chemistry applications [81].
Polymerism.
Many problems in chemistry and materials science involve polymers. By leveraging lists and recursive data structures, the ADT can be extended to represent polymeric and macromolecular architectures, enabling analysis and optimization of complex polymer systems.
Quantum information (spin).
Incorporating basic quantum descriptors (e.g., spin multiplicity, localization of unpaired electrons) would allow the ADT to capture open-shell species and states relevant to reactivity and spectroscopy.
Empirical evaluation.
With these extensions and additional parsers in place, the ADT can be stress-tested across representative cheminformatics benchmarks to quantify downstream impact and validate its practical utility.
3.10.3 Refinement Types via Liquid Haskell
Indeed, one can also use Liquid Haskell [82] to encode simple forms of type-level constraints, known as refinement types. A refinement type system allows us to refine existing types (e.g., Int, Double, etc.) with logical predicates, thereby adding extra compile-time guarantees about program behavior. Unlike fully general dependent types, refinement types represent a compromise between expressive power and decidability.
By adding refinement types to Haskell, these constraints can be checked at compile time rather than waiting until runtime (or even worse, detecting them only after a failed simulation).
3.10.4 Dependent Types
Dependent types allow types to be parameterised by values, so domain invariants can be stated and checked by the type system at compile time. In our setting, they would permit encoding chemical constraints such as valence limits, charge balance, and valid connectivity directly in the type of a molecule, making many physically invalid structures unrepresentable [83]. However, even with such expressive types, it remains unclear how one would ensure that the resulting probabilistic models employ chemically plausible priors or that their likelihood functions correspond to real, experimentally meaningful measurements.
3.10.5 Development of a powerful, user-friendly library
Future additions to the library can use Haskell’s accelerate library [84] to parallelise computations from using monoidal structures (e.g., reductions, scans) being one common pattern for parallelism, accelerate supports a broad range of data-parallel operations on multidimensional arrays (maps, zips, permutations, stencils, etc.). In future engineering work, data-parallel backends (e.g., Accelerate) could be explored for accelerating specific kernels (e.g., traversal, validation) where benchmarks justify it. [85]). Functions should include the ability to convert between the ADT and other common representations, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indices and may admit multiple coordinate realisations. The permutations of atom indices form a group, and these group properties can be leveraged when designing canonicalisation procedures and symmetry-aware comparisons.
3.10.6 Further enhancing the ADT
To further strengthen the type safety and invariance guarantees of the representation, exploring dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [86] could simplify the manipulation of complex molecular structures by providing a composable, modular and type-safe way to access and modify nested data fields.
To enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [87] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics.
These directions would clarify the practical scope of the representation and provide evidence for (or against) its utility on standard cheminformatics tasks, such as accelerating the discovery of novel and useful molecules. Implementing backpropagation using differentiable programming [88, 89] would enable using gradient-based inference algorithms.
We do not consider quantum-computing approaches here; our focus is the representational substrate and its validation properties.
3.10.7 Experimental validation of the ADT
At present, the efficiency and efficacy of the ADT in cheminformatics tasks are hypotheses. To validate the utility of the ADT as a molecular representation, empirical work is needed to benchmark its performance against representations such as SMILES and SELFIES, on tasks like molecular property prediction (e.g. calculation of logP and logS), virtual screening, and de novo drug design.
Another avenue for future work is using the proposed representation with geometric deep learning techniques to exploit symmetries and invariances in molecular structures [2]. By leveraging mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through type classes, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the choice of index for each atomID used in the representation should not influence a model, given the same molecule but with a different ordering of atomic IDs.
An additional avenue is using equational reasoning to investigate algebraic approaches in molecular modeling [90].
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models, but this would have to be demonstrated.
4 Conclusion
In this work, we introduced a molecular representation grounded in algebraic data types and provided a reference implementation to make the method concrete and reproducible. The central idea is to treat molecules as structured, typed values—rather than as strings—so that validation, transformation, and composition are defined directly over the molecular structure and many malformed manipulations can be rejected early.
Our core constitution layer follows a Dietz-style valence multigraph formulation, which supports delocalised and multicentre bonding in a uniform way. We pair this with an 3D configuration layer (coordinates and stereochemical distinctions) and sketch how electronic annotations (shells, subshells, orbitals) can be carried when such metadata are available. We also show how the same typed approach naturally extends to reaction representations, where conservation checks and structural constraints can be expressed as ordinary program logic.
Overall, the contribution is a representation and reference implementation intended to support clearer reasoning, safer manipulation, and tighter integration with Bayesian and geometric machine-learning workflows than is typical with string-first formats.
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we explored group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
The ADT presented was not intended to be a definitive representation, but a valuable representational concept for cheminformaticians.Nevertheless, the ADT provides a structured alternative to string-first formats (e.g., SMILES/SELFIES) for representing, transforming, and validating molecular structures, particularly in workflows that benefit from explicit invariants and compositional operations.
5 Declarations
5.1 Availability and requirements
Project name:
Project name: MolADT-Bayes (reference implementation for this paper)
Project home page:
Project home page: [91]
Archived version:
Operating system(s):
Tested on macOS Tahoe 26.2 (Build 24F74) (GHC 9.6.5, Cabal 3.14.2.0, Stack 3.7.1, ghcup 0.1.50.2)
Programming language:
Haskell (GHC; Stack or Cabal)
Other requirements:
GHC via Stack; build with stack build; run stack exec moladtbayes. A step-by-step README.md documents OS, dependencies, and exact commands.
License:
AGPL-3.0 (OSI-approved)
Restrictions to use by non-academics:
None beyond AGPL-3.0 obligations
The DB1/DB2 naming convention on the GitHub follows the dataset definitions introduced by Donyapour et al. in their ClassicalGSG study of the SAMPL7 logP challenge [92]. In that work, the authors construct a master training corpus (denoted DB1) by aggregating several publicly available experimental logP datasets (summarised in their Table 1). A chemically restricted subset, DB2, is then obtained by filtering DB1 to molecules containing only the elements C, N, O, S, and H, in order to match the elemental composition of the SAMPL7 target compounds. The same DB1/DB2 split is mirrored in our distribution (MolADT-Bayes/logp/DB1.sdf and MolADT-Bayes/logp/DB2.sdf) to enable direct comparability with prior results. These datasets, together with multiple predefined training/test splits and associated molecular and property files, are publicly available as the “ClassicalGSG logP dataset” via Zenodo [72].
Use of AI-assisted editing
We used OpenAI’s GPT-5 Pro (ChatGPT) only for light editorial assistance (copy-editing, LaTeX hygiene, float placement). All domain content—mathematics, chemistry, Haskell code, experiments, and conclusions—was written by the authors; some code refactoring drew on OpenAI’s Codex. Any LLM suggestions were edited, verified, and cited. No proprietary or human-subject data were provided. The authors take full responsibility for the manuscript and codebase.
5.2 Competing interests
The authors declare no competing interests. The repository is already listed on an open source molecular modeling index on GitHub [93].
5.3 Funding
Oliver Goldstein was funded in part by an EPSRC DTP Scholarship and is also partly self-funded.
Sam March is self-funded.
5.4 Authors’ contributions
5.4.1 Oliver Goldstein
Ideated the project and conceptualized the solution in the form of an Algebraic Data Type, implemented the reference prototype and the worked examples/experiments. Oliver drafted the initial article and finalized its writing.
5.4.2 Samuel March
Made substantial revisions to the article’s form and content, contributed to the research, and made miscellaneous contributions to the library.
5.5 Acknowledgments
We are thankful for discussions at the Algebra of Programming group at Oxford University, and comments by Tom Smeding, Sean Moss and Swaraj Dash.
6 Appendices
1benzenePretty :: Molecule
2benzenePretty = Molecule
3 { atoms = atomTable
4 , localBonds = sigmaFramework
5 , systems = [(SystemId 1, piRingSystem)]
6 }
7 where
8 -- Atom IDs
9 carbons = AtomId <\$> [1..6]
10 hydrogens = AtomId <\$> [7..12]
11
12 -- Shared element data (kept shared, like your original)
13 carbonAttributes = elementAttributes C
14 hydrogenAttributes = elementAttributes H
15 carbonShells = elementShells C
16 hydrogenShells = elementShells H
17
18 -- Small helpers
19 coord (x,y,z) =
20 Coordinate (mkAngstrom x) (mkAngstrom y) (mkAngstrom z)
21
22 mkAtom aid attrs sh xyz = Atom
23 { atomID = aid
24 , attributes = attrs
25 , coordinate = coord xyz
26 , shells = sh
27 , formalCharge = 0
28 }
29
30 mkEdges = S.fromList . map (uncurry mkEdge)
31
32 -- Geometry (same numbers, just data-driven)
33 carbonCoords =
34 [ (-1.2131, -0.6884, 0.0)
35 , (-1.2028, 0.7064, 0.0)
36 , (-0.0103, -1.3948, 0.0)
37 , ( 0.0104, 1.3948, 0.0)
38 , ( 1.2028, -0.7063, 0.0)
39 , ( 1.2131, 0.6884, 0.0)
40 ]
41
42 hydrogenCoords =
43 [ (-2.1577, -1.2244, 0.0)
44 , (-2.1393, 1.2564, 0.0)
45 , (-0.0184, -2.4809, 0.0)
46 , ( 0.0184, 2.4808, 0.0)
47 , ( 2.1394, -1.2563, 0.0)
48 , ( 2.1577, 1.2245, 0.0)
49 ]
50
51 carbonAtoms =
[…]
53 carbons
54 carbonCoords
55
56 hydrogenAtoms =
[…]
58 hydrogens
59 hydrogenCoords
60
61 allAtoms = carbonAtoms ++ hydrogenAtoms
62
63 atomTable = M.fromList [(atomID a, a) | a <- allAtoms]
64
65 -- Bonds
66 ringPairs = zip carbons (tail (cycle carbons)) -- (c1,c2) ... (c6,c1)
67 chPairs = zip carbons hydrogens -- (c1,h7) ... (c6,h12)
68
69 sigmaFramework = mkEdges (ringPairs ++ chPairs)
70 piRingEdges = mkEdges ringPairs
71
72 piRingSystem = mkBondingSystem (NonNegative 6) piRingEdges (Just "pi_ring")
Listing A.1: Benzene with one Dietz $\pi$ pool ($s=6$) over the ring edges plus $\sigma$ edges; coordinates in Å. Source : [94].